OpenStudy (anonymous):

Bobby is bringing his collection of dimes and quarters to the bank to deposit into his savings account. If d represents the number of dimes in Bobby’s collection, q represents the number of quarters, and the collection is worth $63.75, which of the following equations models the value of Bobby’s collection?

OpenStudy (anonymous):

A. 63.75 = d + q B. 63.75 = 25d + 10q C. 63.75 = 10d + 25q D. 63.75 = 0.10d + 0.25q

OpenStudy (anonymous):

@BassCatcher15 @dumbcow @mathmate @shamil98 @zepdrix

OpenStudy (anonymous):

A. Since we dont have knowledge of specific amounts of each dime/quarter, we can just say d + q.

OpenStudy (anonymous):

OK thanks dude. Hope you catch more fish next time you go fishing :)

OpenStudy (anonymous):

Haha! Thanks and no problem. Tag me if you need any more help.

OpenStudy (anonymous):

k